3. Using Laplace transform, solve the differential equation y +2y +y=te* given that y(0) = 1, y(0)= -2.
Answer #1

y tay + y = tet yo! y(o)=– 2 la place on both sides Lly + 3y ty] het [Śyos) -sy@_y ©]+2{sys) -YO)449) (6+12 y = 348-sy@3-St1-1 + ysia (Stigt (5+12 et Y(S) = ㅗ stl (5+1) (Stigt inresse laplace on both sides LgSD = 1 [sti (8+12 + shit is that ) = ?